Operations Report – February 4th

Crew 204 Operations Report 04- FEB – 2019

SOL: 7

Name of person filing report: Kunal Naik (Space_bot)

Non-nominal systems: None

Notes on non-nominal systems: N/A

Generator (hours run): Turned on this evening by 1730 (4th Feb),
Stopped by this morning 0800 (4th Feb)

Solar— SOC % – Turned on (4th Feb) 100%, Turned off (4th Feb) 76%

Diesel Reading – 38%

Propane Reading – ~ 65%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 70%, ~ 350 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 219 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01407956 units

Toilet tank emptied: Yes

Deimos rover used: No, still not functional
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Not Applicable
Beginning charge: Not Applicable
Ending charge: Not Applicable
Currently charging: Not Applicable

Spirit rover used: Not Used
Hours: 72.1
Beginning charge: Not Used
Ending charge: Not Used
Currently charging: Yes

Opportunity rover used: Not Used
Hours: 47.6 hours
Beginning charge: Not Used
Ending charge: Not Used
Currently charging: Yes

Curiosity rover used: Not Used
Hours: Not Used
Beginning charge: Not Used
Ending charge: Not Used
Currently charging: Not Used

Notes on rovers: None

ATV’s Used: None

Reason for use: N/A

Oil Added? N/A

ATV Fuel Used: Nope

# Hours the ATVs were Used today: N/A

Notes on ATVs: N/A

HabCar used and why, where? Used to go to the town

Crew Car used and why, where? Off Campus

General notes and comments: None

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAMM operations: Used some engineering tools for Rover.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: None

Journalist Report – February 4th

Crew 204 Journalist Report 04 FEB 2019.

Title: Are We Alone?

Author: Sonal Baberwal.

Good morning from Mars. Last night has been a noisy night as we could sense the winds trying to say us something by knocking the doors throughout the night! The feel of isolation was at its best as all we could hear was the noise of the wind which leaves us with a question, are we alone out here in the middle of nowhere? Perhaps nature won’t ever leave you alone. Another day started with looking out from the window and observing the beautiful view of the clouds and the mountains.

The day continues with yummy brunch cooked by Kunal Naik and Avishek Ghosh. We were actually discussing how much we are missing the EVAs due to unsupported weather conditions. The day had been engaged in working on the usual routines where our GreenHab officer had been working on regular care of the GreenHab.

Kunal Naik and Avishek Ghosh had been working on their projects related to the rover and the 3D printing. The day had been an ordinary day just like the others, but the best part of this ordinary day was the extra-ordinary outputs everyone had received.

Very firstly, heartiest congratulations to our baby plant from the sprouts that have grown one day old and he is observing Mars with us! All the fruitful efforts and care of our GreenHab Officer Patil (Cosmos) have turned to be worthy. The new soil compositions prepared by our commander looks promising with some reasonable or unexpected results.

The rover is finally assembled where it can crawl on the Martian terrains and capture its unforgettable moments. The commander has prepared the clay mixture which is ready to be molded into pots from this soil.

The day had been really interesting in terms of the activities we have been performing. Looking forward to a new sunrise with a hope in better weather conditions to support the EVA. This is Sonal Baberwal, HSO and Crew Journalist reporting live from the upper deck of the Habitat! Stay connected to know more about the application of our results.

Greenhab Report – February 4th

Crew 204 Green Hab Report – 4-FEB-2019

GreenHab Officer: Patil (Cosmos)

Environmental control: Heating and Cooling w/ ambient air (5 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 18.3°C

High: 24.2°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 12 gallons

Water in Blue Tank: ~ 219 gallons

Time(s) of watering for crops:

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)
[AT TIME OF WATERING]

1) 09.00 AM 18.3° C / 36 %
2) 12.00 PM 20.6° C / 33 %
3) 15:00 PM 22.3° C / 38 %
4) 18:00 PM 19.1° C / 62 %

Changes to crops: None

Narrative: Today was the day of "the hub garden germination". We were observing planted sprouts in a garden soil tray and other composites of soils. Seed coat has ruptured, root has grown in ground and crop has started taking its shape. We could see good growth in sprouts planted in a tray, and promising growth in sprouts planted in various composites of soil.

Harvest: None

Support/supplies needed: None

Sol Summary – February 4th

Crew 204 Sol Summary Report 04- FEB – 2019

Sol: 7

Summary Title: The Herb Garden Germination

Author’s name: Kunal Naik (Space_bot)

Mission Status: Nominal

Sol Activity Summary: The most selfless living things are the plants. They just believe in giving essentials. For any interplanetary travel or living on other planet. They are the critical part of the life support system. They provide food, water, and oxygen, which are one of the essentials to sustain our self. Maintaining the Green Hab is a big task and an important one.

The crew woke up little late as EVA was cancelled due to strong winds. The roaring gust, the strong winds striking the hab is the only sound the crew is listening from last day on repeat mode. Sipping of Hot Chocolate and discussion over Veganism, MDRS Green Hab, Growing of Herbs via Aquaponic system escalated by an hour. The crew had brunch on Monday morning which is quiet a strange thing but its Mars! Anything, Anytime. The lunch was cooked earlier by Space Bot and Carbon flight as per the rotation while the other members cleaned the kitchen. The crew is facing some internet issues from last day which can be due to the weather but not a big problem.

Post lunch the crew started to work on their assigned projects. All the rover issues were solved, and the bot is ready to test offside with streamed AV. While Crew Commander, Crew Health and Safety Officer (Mars Bag) which were assisted by the Green Hab Officer (Cosmos) looking after old plants as well observing the new plants which were sowed recently in 8 different soil compositions.

A positive news from the Green Hab, the newly planted seeds are coming to life. The EVA was cancelled due to windy conditions, so its time was utilized for recreational activities. The crew has planned to watch a movie post dinner.

Look Ahead Plan: The 3D clay printer will be tested manually using different soil composition and printing some samples. The work on the child rover will be to test the complete system outdoor during a short EVA.

Finger crossed on weather and servers. The sprouts and mix seeds have already started showing good results. Hoping to observe proper growth.

Anomalies in work: None Encountered

Weather: Cloudy, Strong Winds

Crew Physical Status: Nominal

EVA’s location: Cancelled due to strong winds

Reports to be file: 1) Operations Report 2) Green Hab Report 3)
Journalist Report 4) Photos

Support Requested: None.

Science Report – February 03rd

Title: The Martian experimental benefactor.

Author: Avishek Ghosh

1. Processing and preparation of soil

Team: Sonal Baberwal, Avishek Ghosh

Various soil composition (8 variations) have been prepared from MDRS soil, JSC-MARS-1, Charcoal and garden soil stored in Greenhab. Another material is also added with existing soil composition to prevent compaction. Water has been poured into all the prepared soil samples and left overnight. The complex compositions have soaked water quickly even got dried till morning. Other basic compositions were holding the moisture.

2. Experimental Greenhab preparation

Team: Pranit Patil, Sonal Baberwal, Avishek Ghosh

A few sprouted seeds along with some grown-up mustard plants have been transferred and planted to the new 8 different soil compositions that are prepared yesterday. Observation is required on the surviving capacity and growth of the seeds and plants on new soils. A regular growth has also been observed on the sprouted seeds that already planted in regular garden soil.

3. Rover assembly and integration

Kunal Naik (Space_Bot)

The rover has experienced some issues related to power system and wirings which was resolved at the earliest. The battery test and camera performance test have been executed inside RAMM (Engineering lab). It is ready to be operated outdoor during the EVA’s through a newly developed local server connectivity. But the real communication system is yet to be tested.

4. 3D Printer assembly and test run

Avishek Ghosh (Carbon flight)

Today, finally Commander was able to start the printer system and successfully extrude materials through the tube with automatic/manual settings in engineering to develop. It took a bit of time to develop the mechanical assembly but the entire machine responds well during operations. Printed a sample (MARS) of a few layers during the test run. Further upgradation is required to develop complex printed structures.

Journalist Report – January 03rd

Title: The total detox

Author: Sonal Baberwal

Good Morning from Mars!

Have you ever spent a day without internet? If yes how was your experience?

Today had been an awesome weather! All we could hear is winds! Feels like we are all alone on Mars as the silence was roaring with the sound of he winds. It has been a cloudy day today with a shower of rains!

Today was a long day as we did not have an EVA due to unsupported weather conditions! The morning started with breakfast and the crew had been working on various things. The crew initially cleaned the upper deck and cooked lunch. This was followed by work in REM and GreenHab. Where Patil (Cosmos) had been taking care of the plants and showering them the daily necessary water as per scheduled! Kunal and Avishek had been working in the REM on the 3D printer setup and assembling and testing of the child rover.

Today we had a detox day without any availability of internet due to weather conditions. I could see sleeping beauties all around (including me) as we had no EVA and internet! This was followed by some games and watching movie which was really fun! Sometimes being away from internet is really healthy for our inner peace.

I feel blessed to be a part of this crew as this crew is not about just being professional. The crew is made up of people who take care of each other and spend quality time together enjoying every moment! Whether it might be working together in lab, or planting new seeds in GreenHab or having food together or getting prepared for the EVA. I believe every moment blessed one! Life is truly about enjoying the present moment with the people who deserve your presence!

This is Sonal Baberwal, HSO and Crew journalist reporting live from the upper deck of the Hab. Looking forward for the next day with lot of hope and enthusiasm.

Operations Report – February 03rd

SOL: 6

Name of person filing report: Kunal Naik (Space_bot)

Non-nominal systems: None

Notes on non-nominal systems: N/A

Generator (hours run): Turned on this evening by 1730 (3rd Feb), Stopped by this morning 0830 (3rd Feb)

Solar— SOC % – Turned on (3rd Feb) 73%, Turned off (3rd Feb) 100%;

Diesel Reading – 40%

Propane Reading – ~ 68%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 75%, ~ 375 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 231 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01407695 units

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Spirit rover used: Not Used

Hours: 72.1

Beginning charge: Not Used

Ending charge: Not Used

Currently charging: Yes

Opportunity rover used: Not Used

Hours: 47.6 hours

Beginning charge: Not Used

Ending charge: Not Used

Currently charging: Yes

Curiosity rover used: Not Used

Hours: Not Used

Beginning charge: Not Used

Ending charge: Not Used

Currently charging: Not Used

Notes on rovers: None

ATV’s Used: None

Reason for use: N/A

Oil Added? N/A

ATV Fuel Used: Nope

# Hours the ATVs were Used today: N/A

Notes on ATVs: N/A

HabCar used and why, where? Not Used

Crew Car used and why, where? Off Campus

General notes and comments: None

Summary of internet: Not working after 12

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAMM operations: Used some engineering tools for Rover and 3D clay printer project.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: None

GreenHab Report – February 03rd

[category GreenHab-report]

GreenHab Officer: Patil (Cosmos)

Environmental control: Heating and Cooling w/ ambient air (5 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 18.5°C

High: 24.7°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 12 gallons

Water in Blue Tank: ~ 231 gallons

Time(s) of watering for crops:

TIME NOTED TEMPERATURE (° C) / HUMIDITY (%)

[AT TIME OF
WATERING]

1) 08.34 AM 19.6° C / 48 %

2) 12.00 PM 21.7° C / 42 %

3) 15:05 PM 23.6° C / 28 %

4) 18:00 PM 19.4° C / 43 %

Changes to crops: None

Narrative: Today we performed an experiment in GreenHab. We planted sprouts in various composites of soils, including Garden soil, MDRS soil, JSC Mars one soil and charcoal. We are observing every possible change in them.

Harvest: None

Support/supplies needed: None

Sol Summary – February 03rd

Sol: 6

Summary Title: The Weather Complexity

Author’s name: Kunal Naik (Space_bot)

Mission Status: Nominal

Sol Activity Summary: Nature itself is a super power no one can or will be able to control it. Mission Support! We have a problem, Strong winds incoming! The EVA was proposed to Mission Support last night but after their guidance it was cancelled. The crew woke up little late than decided as it was Sunday, started the morning with warm cup of soup and salty crackers with a round table discussion over the usage of water on Earth, Mars and ISS. Post breakfast the crew engineer (Space Bot) started to work on his child rover’s project. The rover had some serious problems with its circuitry which was found out this morning and was resolved at the earliest. The battery check and camera test (A/V) was done inside the RAMM. While Crew Commander (Carbon Flight) was busy working on his 3D clay printer. The printer was integrated with its extruder as a clay syringe. The printer was mounted on the wooden log to support the system. Delicious Sunday lunch was cooked by Crew Health and Safety Officer (Mars Bag) which was assisted by the Green Hab Officer (Cosmos). He was into his daily assigned Green Hab activities looking after old plants as well observing the new plants which were sowed recently in 8 different soil compositions. Post lunch, the crew spent some time watching movie and relaxing their noon. The EVA was cancelled due to windy conditions, so its time was utilized for recreational activities. The crew has planned to play some card and board games post Sunday dinner.

Look Ahead Plan: The 3D clay printer will be tested using a simulant soil and later with different composition, proceeding printing manually on usual basis. The work on the child rover will be to test the complete system outdoor. Finger crossed on weather and servers. The child-rover will be operated for future EVA’s. The sprouts and mix seeds are already planted in the GreenHab, hoping to get good results.

Anomalies in work: None Encountered

Weather: Sunny, Strong Winds

Crew Physical Status: Nominal

EVA’s location: Cancelled due to strong winds

Reports to be filed: 1) Operations Report 2) Green Hab Report 3) Journalist Report 4) Science Report 5) Photos